The bit was raised to a depth of 4185 m for spudding Hole 1040B. This time the barrel did fully stroke and only required 15 kips to extract. It recovered 5.51 m of core material establishing the mudline. Core 170-1040B-2H did not fully stroke but 8.5 m were recovered, and it was decided to attempt a third APC core. However, while attempting to drill out the rathole from Core 170 1040B-2H it became apparent that the hole could not be advanced without a core barrel in place. The APC barrel was recovered and an XCB core barrel was dropped at that point.
Coring with the XCB continued at a slow pace, because the drill collars making up the BHA were well above the seabed. Low bit weight and top drive rpm were used initially and each were increased as penetration advanced into the firm seabed. Even with the drill collars buried below the seabed the ROP was slow. Bit balling was suspected as the primary reason, because the material being cored was not hard enough to dictate such a slow ROP. Coring continued to Core 170 1040B-22X at 4379.2 mbrf when an XCB space sub failed leaving the XCB hard formation cutting shoe in the hole. The failure terminated coring in Hole 1040B at a total depth of 190.2 mbsf.
Head space methane levels varied from 44,529 to 53,995 ppm whereas ethane levels were 7 to 209 ppm. Propane varied from 0 to 6 ppm. Vacutainer data indicated over 100% methane in several cores. Core 170-1040B-16X was under high pressure when recovered, and core started being ejected out of the barrel at the rig floor immediately upon recovery. Standard precautions regarding safety were taken on the rig floor and core receiving platform.
A total of four DVTP measurements was taken beginning after Core 170-1040B-6X and continued on 050 m intervals during XCB coring. The first DVTP measurement was bad.
Several mud sweeps were pumped in the lower 100 mbsf after fill (1.5 to 12.0 m) was identified between connections. Once coring ceased, the hole was filled with 10.5 lb/gal mud and the drill string was tripped out of the hole. While pulling pipe, an overpull of 30 kips was taken between 120.1 and 177.7 mbsf. The drill string cleared the seafloor at 0630 hr 8 November 1996.
Operation at Site 1039 (CR-1)
During the pipe trip, the vessel was offset 1.2 nmi back to Site 1039 to RCB drill Hole 1039C (see Site 1039 section). After completing Hole 1039C at 0115 hr on 11 November, the pipe was secured, placing the bit at 3988.1 m, for the DP transit back to Site 1040 to RCB spud Hole 1040C.
After spacing out the drill pipe to 4148.0 mbrf, the WSTP was deployed and a bottom-water sample was collected. The previously collected sample at this site was not good. A center bit assembly was then pumped downhole and the drill string was lowered to 4189.0 m. Hole 1040C was spudded at 0915 hr on 11 November 1996 and drilling continued to a depth of 4348.3 m (159.3 mbsf) where RCB coring began. An overlap of approximately 30 m was desired with the previously drilled XCB hole at this site. Coring with the RCB began with Core 170-1040C-1R and continued until a hard layer near the décollement was contacted at the bottom of Core 170-1040C 19R at a depth of 342.2 mbsf. Cores 170-1040C-20R and part of 21R were considerably fractured and assumed to be in the décollement zone. Coring continued on from that point with the penetration rate continually increasing until it stabilized around 60-70 m/hr (8-9 minutes rotating time). The rapid ROP was attributed to coring through the underconsolidated sediments beneath the décollement. A gabbro sill was contacted in Core 170-1040C-52R at a depth of 657.0 mbsf and the last core recovered was 170-1040C-53R at a total depth of 665.0 mbsf.
Some drill string drag was identified while drilling Hole 1040C. The Tensor electronic multishot tool was deployed and drift measurements were taken while retrieving Core 170-1040C-43R. Measurements were taken on bottom at 4750 m (573.1 mbsf) and 100 m above bottom at 4650 m (473.1 mbsf). The hole angle was 18.5° and 17.5°, respectively. It is assumed that the hard seafloor, coupled with a noncentralized coring assembly, contributed to the build-up of angle. A complete drift survey was conducted while recovering Core 170-1040C-53R. The data indicated a dog leg at approximately 440 mbsf. Hole deviation continued to build from 8.25° at 40 mbsf to 18.5° at 440 mbsf and then declined to 14° at 640 mbsf.
A total of 350 bbls of Sepiolite drilling mud was circulated during the coring operation, including spotting mud around the drill collars during DVTP temperature measurements. A bottom-water temperature was taken with the Adara tool during the WSTP deployment and then a total of six DVTP measurements were taken. These began after Core 170-1040C-7R and continued in 050 m intervals during RCB coring. The last DVTP measurement was taken after Core 170-1040C-32R, because it was deemed unwise to risk the drill string and hole by sitting on bottom without circulating.
Above the décollement (Cores 170-1040C-1R to 19R), head space methane levels varied from 4818 to 116,556 ppm, whereas ethane levels were 5 to 65 ppm. Propane varied from 3 to 24 ppm. Vacutainer data indicated 50%-100% methane, whereas C2 ranged from 226 to 1561 ppm and C3 was 0 to 169 ppm. Heavier hydrocarbons like IC4 ranged from 0 to 26 ppm and IC5 was 0 to 9 ppm. There was geochemical evidence of hydrate in situ above the décollement, however, no hydrate was recovered in the cores. The hydrate stability zone for this site extended all the way to basement.
Below the décollement (Cores 170-1040C-20R to 51R), head space methane levels dropped from 12,724 to 4 ppm whereas ethane levels went from 15 to 0 ppm. Propane varied from 19 to 0 ppm, with nothing over 1 ppm after Core 170-1040C-21R. The only vacutainer data recovered below the décollement was for Core 170-1040C-21R. Methane indicated 98% while C2 was 781 ppm and C3 was 76 ppm. Heavier hydrocarbons like IC4 and IC5 registered 29 ppm and 1 ppm, respectively. After that there was not enough gas to warrant vacutainer analysis. There was no geochemical evidence of hydrate in situ below the décollement.
Because there was interest in getting wireline sonic and Formation MicroScanner (FMS) logs (not available through LWD), we decided that if the hole remained open another wiper trip would be made followed by two logging runs . A wiper trip was completed to 4278.0 m (89.0 mbsf) noting a maximum overpull and drag of 40 kips. The bit was released and the mechanical bit release (MBR) sleeve was reverse shifted to close off the dog ports. The hole was displaced with 220 bbls of Sepiolite mud and the open ended pipe was positioned for logging at 4292.8 m (103.8 mbsf).
The first suite of logs (sonic) was only able to advance to 129.2 m below the bit before being halted by a bridge or ledge in the hole. This interval was logged and the tools were recovered ending the wireline logging attempt. The open ended pipe was run to a depth of 4480.1 m (291.1 mbsf) where the drag reached 35 kips. The trip was halted at that point and the hole was displaced with 80 bbls of 10.5 lb/gal mud. The drill pipe was pulled out of the hole and cleared the seafloor at 0400 hr. The ship was then DP offset back to Site 1039 while the pipe trip continued to the surface. The MBR reached the rig floor at 1200 hr on 18 November 1996, ending Hole 1040C.
Operations at Site 1039 (CR-1)
The ship returned to Site 1039 to drill the LWD Hole 1039D, which was spudded at 2330 hr on 18 November 1996. An operational summary is given in the Site 1039 section. Hole 1039D was ended at 0915 on 20 November.
The LWD tools from the previous hole were removed from the rig floor and new ones with fresh batteries were madeup while the ship was moved in DP mode to Site 1040. Once the nuclear sources were loaded, the CDN and CDR tools were run to bottom. A Smith tricone F4 bit was run with 12/32" and 14/32" jets installed to allow coring of a limited amount of igneous rock. LWD Hole 1040D was spudded at 1830 hr 20 November 1996. Because of the very hard seafloor, very controlled drilling parameters were used. Weight-on-bit was limited to 2-4 kips and the rotary speed was 30-45 rpm. Penetration rate for the first 40 m was 7.9 m/hr. Drilling proceeded relatively well, although the penetration rate (average 14.6 m/hr) was hampered by continual bit balling and suspected balling around the LWD stabilizer. At a depth of 279.8 mbsf, pump pressure and torque increased and generous Sepiolite mud sweeps did little to correct the condition. Drilling continued to a depth of 308.7 mbsf while hole conditions continued to deteriorate. A wiper trip was made to 103.6 mbsf with overpulls of 20 kips and 300-500 psi back pressure prevalent to 214.0 mbsf. On the way back to bottom there were 10-15 kips of pipe drag in the hole. In all, a total of 10-1/2 hr of remedial work was attempted to stabilize the hole. Drilling resumed and the hole was deepened to a depth of 4526.5 m (337.5 mbsf). At each connection, we took care to double-ream the hole. Continued hole trouble eventually resulted in stuck pipe, which forced a halt to drilling operations at a total depth of 337.5 mbsf. An overpull of 50 kips coupled with 800 amps torque was required to free the drill string.
Suspected hole deviation and possible formation overpressure was suspected as the primary cause for failure to achieve the depth objective. After pulling pipe to a "safe" depth of 4475.0 m (286.0 mbsf), the Tensor electronic multishot was deployed for a drift survey of the hole. Because of poor hole conditions and the nuclear sources installed in the LWD tools, the lowest data point recorded was at 271.0 mbsf, which was 66.5 meters above hole TD. Drift data indicated that the hole was started at an angle of 7.6° from vertical and proceeded to straighten with depth. The angle was 2.9° at a depth of 271.0 mbsf. After completing the survey, the pipe stuck again. This time 25 kips overpull and 700 amps torque were required to free the string. The hole was displaced with 10.5 lb/gal mud and the drill string tripped back to the surface. Once the nuclear sources were removed from the LWD tools, the bit was pulled to the rig floor. The log data were transferred from the tools to a workstation computer for plotting and analysis. Subsequent analysis of the LWD log data by the logging engineer confirmed a probable overpressured zone in the hole extending from a depth of 281.0 mbsf to TD.
Hole 1040E was located 30 m north of Hole 1040D. We made another attempt to get LWD logs down to and across the décollement. The CDN and CDR tools were changed out for those with fresh batteries, and the same Smith FDGH tricone drag bit that was used successfully in Hole 1039D was run. We hoped that a faster penetration rate and a straighter hole would allow this hole to reach a deeper TD. Drilling in this hole proceeded at a much faster ROP with this more efficient drag-type bit. The first 221.9 m were drilled at an average ROP of 30.6 m/hr. Drilling continued without incident until a depth of around 280 mbsf. At that point hole problems similar to those encountered in Hole 1040D began. Pump pressure (400-500 psi back pressure) and drilling torque (400-500 amps) were higher and tight hole (20-30 kips) conditions were the rule rather than the exception. The hole was advanced to a depth of 4507.4 m (318.4 mbsf) when the drill string began sticking every time a connection was made and it was apparent that the hole could not be drilled much farther. The hole was terminated to avoid the potential loss of the expensive LWD tools. The drill string was raised to 247.9 mbsf and the hole was displaced with 10.5 lb/gal mud. The pipe was pulled clear of the seafloor at 1130 hr. The vessel immediately got underway in DP mode for Site 1041 while the LWD tools were removed from the rig floor and the APC/XCB BHA for the next site was made up.
To 170 Operations - Site 1041
To 170 Table of Contents